# BevSplat: Resolving Height Ambiguity via Feature-Based Gaussian Primitives for Weakly-Supervised Cross-View Localization

### v3

@@ Abstract
This paper addresses the problem of weakly supervised cross-view localization, where the goal is to estimate the pose of a ground camera relative to a satellite image with noisy ground truth annotations. A common approach to bridge the cross-view domain gap for pose estimation is Bird’s-Eye View (BEV) synthesis. However, existing methods struggle with height ambiguity due to the lack of depth information in ground images and satellite height maps. Because a single 2D pixel could represent points at various depths and heights, its true 3D position is ambiguous. Previous solutions either assume a flat ground plane or rely on complex models, such as cross-view transformers. We propose BevSplat, a novel method that resolves height ambiguity by using feature-based Gaussian primitives. Each pixel in the ground image is represented by a 3D Gaussian with semantic and spatial features, which are synthesized into a BEV feature map for relative pose estimation. We validate our method on the widely used KITTI and VIGOR datasets, which include both pinhole and panoramic query images. Experimental results show that BevSplat significantly improves localization accuracy over prior approaches.
@@ 1 Inotroduction
To navigate these complexities, particularly the common difficulty of acquiring precise ground-truth (GT) camera locations at scale, weakly supervised learning [1, 2] has recently emerged as a promising paradigm. In this setting, models are trained using only noisy annotations, such as approximate camera locations with errors potentially reaching tens of meters, which adds another layer of complexity to the task. Nevertheless, the primary advantage of weak supervision lies in its ability to leverage less labor-intensive data collection, making it a more scalable and practical approach for many real-world applications where extensive precise annotations are infeasible.
A key strategy to address cross-view localization is Bird’s-Eye View (BEV) synthesis [3, 4, 5, 1, 6], which generates a bird’s-eye view representation from the ground-level image. The BEV image can then be compared directly to a satellite image, facilitating relative pose estimation. However, existing methods often rely on Inverse Perspective Mapping (IPM), which assumes a flat ground plane [1, 6], or on high-complexity models like cross-view transformers [3, 4, 5] to address height ambiguity, the challenge of resolving the elevation difference between the ground and satellite views.
The flat terrain assumption used in IPM leads to the loss of critical scene information above the ground plane and introduces distortions for objects farther from the camera, as shown in Fig. 1. On the other hand, while cross-view transformers are effective at handling distortions and objects above the ground plane, they are computationally expensive. Furthermore, in weakly supervised settings, noisy ground camera pose annotations provide weak supervision, making it difficult for high-complexity models like transformers to converge [1], ultimately leading to suboptimal localization performance.
In this paper, we propose BevSplat to address these challenges. BevSplat generates feature-based 3D Gaussian primitives for BEV synthesis. Unlike previous 3D Gaussian Splatting (3DGS) [7] methods that rely on color-based representations, we represent each pixel in the ground-level image as a 3D Gaussian with semantic and spatial features. These Gaussians are associated with attributes such as position in 3D space, scale, rotation, and density, which are synthesized into a BEV feature map using a visibility-aware rendering algorithm that supports anisotropic splatting. This approach enables us to handle height ambiguity and complex cross-view occlusions, improving the alignment between the ground-level image and the satellite view for more accurate pose estimation, without the need for expensive depth sensors or complex model architectures.
@@ 2.1 Cross-view Localization
Cross-view localization, which aligns ground-level images with satellite imagery, has evolved from image retrieval to fine-grained pose estimation. Early approaches framed this as an image retrieval task, using metric learning to match ground queries to satellite image slices [8, 9, 10, 11, 12]. While modern transformers have improved retrieval performance, practical application remains challenging [13, 14]. Many recent methods adopt a coarse-to-fine pipeline. This typically involves a coarse retrieval step [15] followed by fine-grained (pixel-level) localization to identify the precise camera pose [16, 17, 18, 3, 19, 5, 20, 21, 22]. A key limitation of these methods is their reliance on precise, GPS-based training data, which is often prone to inaccuracies. To overcome this, weakly supervised settings have been proposed to learn from noisy pose annotations [1, 2]. These weakly supervised settings differ in their assumptions. [2] assumes the availability of GT labels in a source domain and access to cross-view pairs in the target domain. In contrast, [1] addresses a more challenging scenario where source domain GT labels are unavailable and no target domain pairs are accessible. In this work, we tackle the same task setting as [1].
@@ 2.2 Bird’s-Eye View Synthesis
BEV synthesis, which generates bird’s-eye view images from ground-level perspectives, has been widely applied to cross-view localization. While LiDAR and Radar sensors offer high accuracy for localization tasks [23, 24, 25, 26], their high cost limits their use. For camera-only systems, multi-camera setups are commonly employed [27, 28, 29, 30], primarily focusing on tasks like segmentation and recognition. In localization, methods like Inverse Perspective Mapping (IMP) assume a flat ground plane for BEV synthesis [1, 6, 20, 18], which can be overly simplistic for complex environments. Transformer-based models address these challenges but struggle with weak supervision and noisy pose annotations [3, 4, 5]. While methods such as [31, 32] also employ feature Gaussians for the image-to-BEV transformation, they typically benefit from rich depth and semantic information afforded by multi-sensor setups. While effective in some contexts, they face limitations in resource-constrained, real-world scenarios. In stark contrast, our approach is constrained to rely exclusively on weakly supervised signals derived purely from images, presenting a considerably more challenging task.
@@ 2.3 Sparse-View 3D Reconstruction
In our method, we adopt algorithms similar to 3D reconstruction to represent ground scenes. Sparse-view 3D reconstruction has been a major focus of the community. Nerf-based approaches [33] and their adaptations [34] have shown the potential for single-view 3D reconstruction, though their application is limited by small-scale scenes and high computational cost. Recent works using diffusion models [35, 36, 37] and 3D Gaussian representations [7, 38, 39, 40], as well as transformer- and Gaussian-based models[41, 42], have achieved sparse-view 3D reconstruction on a larger scale, but the complexity of these models still restricts their use due to computational demands. Approaches like [43, 44, 45] leverage pre-trained models to directly generate Gaussian primitives, avoiding the limitations of complex models while enabling scene reconstruction from sparse views. We apply such methods to single-view reconstruction, achieving high-accuracy cross-view localization.
@@ 3.1 Geometric Gaussian Primitives Generation
Inspired by 3D Gaussian Splatting (3DGS) [7], we represent the 3D scene as a collection of Gaussian primitives. Our generation process first establishes their initial geometry and appearance. Given the inherent difficulty of directly learning accurate depth in our weakly supervised framework, we utilize a pre-trained depth estimation model to predict per-pixel depth $D_{i}$ from the ground-level image. The initial 3D coordinate $\mu_{i}$ for primitives associated with each pixel $(u_{i},v_{i})$ is then determined from $D_{i}$ and the specific camera model.
@@
Subsequently, a ResNet [46] extracts local features $\mathbf{f}_{\text{loc}}^{i}$ for each pixel $i$ from the ground-level image. These features serve as input to a multi-layer perceptron (MLP [47]), denoted $F_{gs}$, which predicts attributes for $N_{p}=3$ distinct Gaussian primitives originating from each pixel. The predicted attributes for each of these $N_{p}$ Gaussian primitives include positional offsets $\Delta\mathbf{p}_{k}=(\Delta x_{k},\Delta y_{k},\Delta z_{k})$ relative to $\mu_{i}$, an anisotropic scale $\mathbf{S}_{k}$, a rotation quaternion $\mathbf{R}_{k}$, and an opacity value $O_{k}$. Predicting multiple primitives per pixel enhances single-image representation density. The collection of parameters $G_{i}$ for these $N_{p}$ primitives from pixel $i$, incorporating their final 3D positions, is:
@@
Figure 2: BevSplat Framework Overview. Query ground image Gaussian primitive initialization involves: (1) A pre-trained depth model for initial 3D positions ($\mu_{i}$). (2) A ResNet and MLP to predict offsets ($\Delta\mathbf{p}_{k}$), scale ($\mathbf{S}_{k}$), rotation ($\mathbf{R}_{k}$), and opacity ($O_{k}$). (3) A DPT-fine-tuned DINOv2 for extracting semantic features ($\mathbf{f}_{i}$) and confidences ($c_{i}$), which are then bound to these Gaussians. These feature Gaussians are subsequently rendered into BEV feature and confidence maps. Satellite image features are extracted using an identical DINOv2-DPT backbone (note: weights are shared for KITTI but differ for VIGOR, similar to G2SWeakly [1]). Localization is achieved by matching satellite features with the rendered query BEV features via cosine similarity within a sliding window.
@@ 3.2 Feature-based Gaussian Primitives for Relative Pose Estimation
For robust semantic feature extraction from ground and satellite images, inspired by [43, 48, 44], we fine-tune a pre-trained DINOv2 [49] model augmented with a DPT [50] module. From the ground image, this pipeline yields a feature map $\mathbf{F}_{g}\in\mathbb{R}^{H_{g}\times W_{g}\times C}$ and a confidence map $\mathbf{C}_{g}\in\mathbb{R}^{H_{g}\times W_{g}\times 1}$. The confidence map $\mathbf{C}_{g}$, derived from $\mathbf{F}_{g}$ via an additional convolutional layer and a sigmoid activation, assigns lower weights to dynamic objects (e.g., vehicles) and higher weights to static elements (e.g., road surfaces), indicating feature reliability for localization. For the predominantly static satellite image, we solely extract its feature map $\mathbf{F}_{s}\in\mathbb{R}^{H_{s}\times W_{s}\times C}$.
@@ 3.2.1 BEV Feature Rendering
Next, viewing the scene from a BEV perspective (camera directed downwards), the features $\mathbf{f}_{b}$ and confidences $c_{b}$ bound to each Gaussian primitive $b$ are rendered onto a 2D plane. This differentiable $\alpha$-blending process, analogous to RGB rendering in 3DGS [7], yields the BEV feature map $\mathbf{F}_{BEV}$ and confidence map $\mathbf{C}_{BEV}$:
@@ 3.2.2 Pose Estimation via Confidence-Guided Similarity Learning
Supervision. Following [1], a deep metric learning objective supervises the network. For each query ground image, we compute location probability maps: $\mathbf{P}_{\text{pos}}$ against its positive satellite image and $\mathbf{P}_{\text{neg},\text{idx}}$ against each of the $M$ negative satellite images. The weakly supervised loss $\mathcal{L}_{\text{Weakly}}$ aims to maximize the peak of $\mathbf{P}_{\text{pos}}$ while minimizing the peak for each $\mathbf{P}_{\text{neg},\text{idx}}$:
@@ 4 Experiments
KITTI dataset. The KITTI dataset [51] consists of ground-level images captured by a forward-facing pinhole camera with a restricted field of view, complemented by aerial images [52], where each aerial patch covers a ground area of approximately $100\times 100\mathrm{m}^{2}$. The dataset includes a training set and two test sets(Same-Area and Cross-Area). For the Same-Area test set, the test query images are from the same geographical regions as the training set, but are not the same images. For the Cross-Area test set, the test images come from entirely new geographical regions that were not seen during training, testing the model’s ability to generalize. The location search range of ground images is approximately $56\times 56\mathrm{m}^{2}$, with an orientation noise of $\pm 10^{\circ}$.
VIGOR dataset. The VIGOR dataset [15] includes geo-tagged ground panoramas and satellite images from four US cities: Chicago, New York, San Francisco, and Seattle. Each satellite patch spans $70\times 70\mathrm{m}^{2}$ and is labeled positive if the ground camera is within its central $1/4$ region; otherwise, it is semi-positive. The dataset also has Same-Area and Cross-Area splits: Same-Area uses training and testing data from the same region, while Cross-Area splits training and testing between two separate city groups. We use only positive satellite images for all experiments, following [1].
@@
Evaluation Metrics. For the KITTI dataset [51], we evaluated localization and orientation errors by calculating mean and median errors in meters and degrees, respectively. We also compute recall at thresholds of 1 m and 3 m for longitudinal (along the driving direction) and lateral (orthogonal to the driving direction) localization errors, as well as 1 ° and 3 ° for orientation errors. A localization is considered successful if the estimated position falls within the threshold of the ground truth, and an orientation is accurate if its error is within the angle threshold. For the VIGOR dataset [15], which does not provide driving direction information, we report mean and median errors as outlined in [1].
@@
Implementation Details. For 3D point cloud generation from ground images, we employ specific depth estimation models: DepthAnythingV2 [53] for the pinhole camera images of the KITTI dataset [51], and UniK3D [54] for the panoramic images of the VIGOR dataset [15]. Our feature extractor for both ground and satellite images is a DINOv2 backbone [49], initialized with FiT weights [48], which is subsequently fine-tuned using an attached DPT module [50]. This extractor yields satellite feature maps with dimensions $(C,H,W)=(32,128,128)$. For ground images, initial feature maps of $(32,64,256)$ are extracted; these are then projected into BEV using our feature Gaussians projection, resulting in final ground BEV features also of dimensions $(32,128,128)$. Our model is trained using the AdamW optimizer [55] (weight decay $10^{-3}$) and a OneCycleLR scheduler [56] with a cosine annealing strategy, where the learning rate peaks at $6.25\times 10^{-5}$. We use a batch size of 8 on a single 4090 GPU, training for 8 epochs on KITTI and 14cf epochs on VIGOR.

### v4

@@ Abstract
This paper addresses the problem of weakly supervised cross-view localization, where the goal is to estimate the pose of a ground camera relative to a satellite image with noisy ground truth annotations. A common approach to bridge the cross-view domain gap for pose estimation is Bird’s-Eye View (BEV) synthesis. However, existing methods struggle with height ambiguity due to the lack of depth information in ground images and satellite height maps. Because a single 2D pixel could represent points at various depths and heights, its true 3D position is ambiguous. Previous solutions either assume a flat ground plane or rely on complex models, such as cross-view transformers. We propose BevSplat, a novel method that resolves height ambiguity by using feature-based Gaussian primitives. Each pixel in the ground image is represented by a 3D Gaussian with semantic and spatial features, which are synthesized into a BEV feature map for relative pose estimation. We validate our method on the widely used KITTI and VIGOR datasets, which include both pinhole and panoramic query images. Experimental results show that BevSplat significantly improves localization accuracy over prior approaches. Our code is available at https://github.com/wangqww/BevSplat.
@@ 1 Inotroduction
To navigate these complexities, particularly the common difficulty of acquiring precise ground-truth (GT) camera locations at scale, weakly supervised learning [shi2024weakly, xia2024adapting] has recently emerged as a promising paradigm. In this setting, models are trained using only noisy annotations, such as approximate camera locations with errors potentially reaching tens of meters, which adds another layer of complexity to the task. Nevertheless, the primary advantage of weak supervision lies in its ability to leverage less labor-intensive data collection, making it a more scalable and practical approach for many real-world applications where extensive precise annotations are infeasible.
A key strategy to address cross-view localization is Bird’s-Eye View (BEV) synthesis [fervers2022uncertainty, shi2023boosting, sarlin2023orienternet, shi2024weakly, wang2024fine], which generates a bird’s-eye view representation from the ground-level image. The BEV image can then be compared directly to a satellite image, facilitating relative pose estimation. However, existing methods often rely on Inverse Perspective Mapping (IPM), which assumes a flat ground plane [shi2024weakly, wang2024fine], or on high-complexity models like cross-view transformers [fervers2022uncertainty, shi2023boosting, sarlin2023orienternet] to address height ambiguity, the challenge of resolving the elevation difference between the ground and satellite views.
The flat terrain assumption used in IPM leads to the loss of critical scene information above the ground plane and introduces distortions for objects farther from the camera, as shown in Fig. 1. On the other hand, while cross-view transformers are effective at handling distortions and objects above the ground plane, they are computationally expensive. Furthermore, in weakly supervised settings, noisy ground camera pose annotations provide weak supervision, making it difficult for high-complexity models like transformers to converge [shi2024weakly], ultimately leading to suboptimal localization performance.
In this paper, we propose BevSplat to address these challenges. BevSplat generates feature-based 3D Gaussian primitives for BEV synthesis. Unlike previous 3D Gaussian Splatting (3DGS) [kerbl20233d] methods that rely on color-based representations, we represent each pixel in the ground-level image as a 3D Gaussian with semantic and spatial features. These Gaussians are associated with attributes such as position in 3D space, scale, rotation, and density, which are synthesized into a BEV feature map using a visibility-aware rendering algorithm that supports anisotropic splatting. This approach enables us to handle height ambiguity and complex cross-view occlusions, improving the alignment between the ground-level image and the satellite view for more accurate pose estimation, without the need for expensive depth sensors or complex model architectures.
@@ 2.1 Cross-view Localization
Cross-view localization, which aligns ground-level images with satellite imagery, has evolved from image retrieval to fine-grained pose estimation. Early approaches framed this as an image retrieval task, using metric learning to match ground queries to satellite image slices [lin2013cross, regmi2018cross, shi2019spatial, Liu_2019_CVPR, shi2020optimal]. While modern transformers have improved retrieval performance, practical application remains challenging [yang2021cross, Zhu_2022_CVPR]. Many recent methods adopt a coarse-to-fine pipeline. This typically involves a coarse retrieval step [zhu2021vigor] followed by fine-grained (pixel-level) localization to identify the precise camera pose [shi2022beyond, xia2022visual, song2024learning, fervers2022uncertainty, lentsch2023slicematch, sarlin2023orienternet, wang2024view, xia2025fg, lee2025pidloc]. A key limitation of these methods is their reliance on precise, GPS-based training data, which is often prone to inaccuracies. To overcome this, weakly supervised settings have been proposed to learn from noisy pose annotations [shi2024weakly, xia2024adapting]. These weakly supervised settings differ in their assumptions. [xia2024adapting] assumes the availability of GT labels in a source domain and access to cross-view pairs in the target domain. In contrast, [shi2024weakly] addresses a more challenging scenario where source domain GT labels are unavailable and no target domain pairs are accessible. In this work, we tackle the same task setting as [shi2024weakly].
@@ 2.2 Bird’s-Eye View Synthesis
BEV synthesis, which generates bird’s-eye view images from ground-level perspectives, has been widely applied to cross-view localization. While LiDAR and Radar sensors offer high accuracy for localization tasks [qin2023supfusion, harley2023simple, lin2024rcbevdet, liu2025seed], their high cost limits their use. For camera-only systems, multi-camera setups are commonly employed [reiher2020sim2real, li2022bevformer, yang2023parametric, sarlin2023snap], primarily focusing on tasks like segmentation and recognition. In localization, methods like Inverse Perspective Mapping (IMP) assume a flat ground plane for BEV synthesis [shi2024weakly, wang2024fine, wang2024view, song2024learning], which can be overly simplistic for complex environments. Transformer-based models address these challenges but struggle with weak supervision and noisy pose annotations [fervers2022uncertainty, shi2023boosting, sarlin2023orienternet]. While methods such as [chabot2025gaussianbev, lu2025gaussianlss] also employ feature Gaussians for the image-to-BEV transformation, they typically benefit from rich depth and semantic information afforded by multi-sensor setups. While effective in some contexts, they face limitations in resource-constrained, real-world scenarios. In stark contrast, our approach is constrained to rely exclusively on weakly supervised signals derived purely from images, presenting a considerably more challenging task.
@@ 2.3 Sparse-View 3D Reconstruction
In our method, we adopt algorithms similar to 3D reconstruction to represent ground scenes. Sparse-view 3D reconstruction has been a major focus of the community. Nerf-based approaches [mildenhall2021nerf] and their adaptations [hong2023lrm] have shown the potential for single-view 3D reconstruction, though their application is limited by small-scale scenes and high computational cost. Recent works using diffusion models [rombach2021highresolution, ze2025satdreamer360multiviewconsistentgenerationgroundlevel, ze2025controllablesatellitetostreetviewsynthesisprecise] and 3D Gaussian representations [kerbl20233d, cai2024baking, zhou2024diffgs, mu2025gsd], as well as transformer- and Gaussian-based models[chen2024splatformer, GaussTR], have achieved sparse-view 3D reconstruction on a larger scale, but the complexity of these models still restricts their use due to computational demands. Approaches like [zhou2024feature, wewer2025latentsplat, szymanowicz2024flash3d] leverage pre-trained models to directly generate Gaussian primitives, avoiding the limitations of complex models while enabling scene reconstruction from sparse views. We apply such methods to single-view reconstruction, achieving high-accuracy cross-view localization.
@@ 3.1 Geometric Gaussian Primitives Generation
Inspired by 3D Gaussian Splatting (3DGS) [kerbl20233d], we represent the 3D scene as a collection of Gaussian primitives. Our generation process first establishes their initial geometry and appearance. Given the inherent difficulty of directly learning accurate depth in our weakly supervised framework, we utilize a pre-trained depth estimation model to predict per-pixel depth $D_{i}$ from the ground-level image. The initial 3D coordinate $\mu_{i}$ for primitives associated with each pixel $(u_{i},v_{i})$ is then determined from $D_{i}$ and the specific camera model.
@@
Subsequently, a ResNet [he2016deep] extracts local features $\mathbf{f}_{\text{loc}}^{i}$ for each pixel $i$ from the ground-level image. These features serve as input to a multi-layer perceptron (MLP [rosenblatt1958perceptron]), denoted $F_{gs}$, which predicts attributes for $N_{p}=3$ distinct Gaussian primitives originating from each pixel. The predicted attributes for each of these $N_{p}$ Gaussian primitives include positional offsets $\Delta\mathbf{p}_{k}=(\Delta x_{k},\Delta y_{k},\Delta z_{k})$ relative to $\mu_{i}$, an anisotropic scale $\mathbf{S}_{k}$, a rotation quaternion $\mathbf{R}_{k}$, and an opacity value $O_{k}$. Predicting multiple primitives per pixel enhances single-image representation density. The collection of parameters $G_{i}$ for these $N_{p}$ primitives from pixel $i$, incorporating their final 3D positions, is:
@@
Figure 2: BevSplat Framework Overview. Query ground image Gaussian primitive initialization involves: (1) A pre-trained depth model for initial 3D positions ($\mu_{i}$). (2) A ResNet and MLP to predict offsets ($\Delta\mathbf{p}_{k}$), scale ($\mathbf{S}_{k}$), rotation ($\mathbf{R}_{k}$), and opacity ($O_{k}$). (3) A DPT-fine-tuned DINOv2 for extracting semantic features ($\mathbf{f}_{i}$) and confidences ($c_{i}$), which are then bound to these Gaussians. These feature Gaussians are subsequently rendered into BEV feature and confidence maps. Satellite image features are extracted using an identical DINOv2-DPT backbone (note: weights are shared for KITTI but differ for VIGOR, similar to G2SWeakly [shi2024weakly]). Localization is achieved by matching satellite features with the rendered query BEV features via cosine similarity within a sliding window.
@@ 3.2 Feature-based Gaussian Primitives for Relative Pose Estimation
For robust semantic feature extraction from ground and satellite images, inspired by [zhou2024feature, yue2025improving, wewer2025latentsplat], we fine-tune a pre-trained DINOv2 [oquab2023dinov2] model augmented with a DPT [ranftl2021vision] module. From the ground image, this pipeline yields a feature map $\mathbf{F}_{g}\in\mathbb{R}^{H_{g}\times W_{g}\times C}$ and a confidence map $\mathbf{C}_{g}\in\mathbb{R}^{H_{g}\times W_{g}\times 1}$. The confidence map $\mathbf{C}_{g}$, derived from $\mathbf{F}_{g}$ via an additional convolutional layer and a sigmoid activation, assigns lower weights to dynamic objects (e.g., vehicles) and higher weights to static elements (e.g., road surfaces), indicating feature reliability for localization. For the predominantly static satellite image, we solely extract its feature map $\mathbf{F}_{s}\in\mathbb{R}^{H_{s}\times W_{s}\times C}$.
@@ 3.2.1 BEV Feature Rendering
Next, viewing the scene from a BEV perspective (camera directed downwards), the features $\mathbf{f}_{b}$ and confidences $c_{b}$ bound to each Gaussian primitive $b$ are rendered onto a 2D plane. This differentiable $\alpha$-blending process, analogous to RGB rendering in 3DGS [kerbl20233d], yields the BEV feature map $\mathbf{F}_{BEV}$ and confidence map $\mathbf{C}_{BEV}$:
@@ 3.2.2 Pose Estimation via Confidence-Guided Similarity Learning
Supervision. Following [shi2024weakly], a deep metric learning objective supervises the network. For each query ground image, we compute location probability maps: $\mathbf{P}_{\text{pos}}$ against its positive satellite image and $\mathbf{P}_{\text{neg},\text{idx}}$ against each of the $M$ negative satellite images. The weakly supervised loss $\mathcal{L}_{\text{Weakly}}$ aims to maximize the peak of $\mathbf{P}_{\text{pos}}$ while minimizing the peak for each $\mathbf{P}_{\text{neg},\text{idx}}$:
@@ 4 Experiments
KITTI dataset. The KITTI dataset [geiger2013vision] consists of ground-level images captured by a forward-facing pinhole camera with a restricted field of view, complemented by aerial images [shi2022accurate], where each aerial patch covers a ground area of approximately $100\times 100\mathrm{m}^{2}$. The dataset includes a training set and two test sets(Same-Area and Cross-Area). For the Same-Area test set, the test query images are from the same geographical regions as the training set, but are not the same images. For the Cross-Area test set, the test images come from entirely new geographical regions that were not seen during training, testing the model’s ability to generalize. The location search range of ground images is approximately $56\times 56\mathrm{m}^{2}$, with an orientation noise of $\pm 10^{\circ}$.
VIGOR dataset. The VIGOR dataset [zhu2021vigor] includes geo-tagged ground panoramas and satellite images from four US cities: Chicago, New York, San Francisco, and Seattle. Each satellite patch spans $70\times 70\mathrm{m}^{2}$ and is labeled positive if the ground camera is within its central $1/4$ region; otherwise, it is semi-positive. The dataset also has Same-Area and Cross-Area splits: Same-Area uses training and testing data from the same region, while Cross-Area splits training and testing between two separate city groups. We use only positive satellite images for all experiments, following [shi2024weakly].
@@
Evaluation Metrics. For the KITTI dataset [geiger2013vision], we evaluated localization and orientation errors by calculating mean and median errors in meters and degrees, respectively. We also compute recall at thresholds of 1 m and 3 m for longitudinal (along the driving direction) and lateral (orthogonal to the driving direction) localization errors, as well as 1 ° and 3 ° for orientation errors. A localization is considered successful if the estimated position falls within the threshold of the ground truth, and an orientation is accurate if its error is within the angle threshold. For the VIGOR dataset [zhu2021vigor], which does not provide driving direction information, we report mean and median errors as outlined in [shi2024weakly].
@@
Implementation Details. For 3D point cloud generation from ground images, we employ specific depth estimation models: DepthAnythingV2 [yang2024depth] for the pinhole camera images of the KITTI dataset [geiger2013vision], and UniK3D [piccinelli2025unik3d] for the panoramic images of the VIGOR dataset [zhu2021vigor]. Our feature extractor for both ground and satellite images is a DINOv2 backbone [oquab2023dinov2], initialized with FiT weights [yue2025improving], which is subsequently fine-tuned using an attached DPT module [ranftl2021vision]. This extractor yields satellite feature maps with dimensions $(C,H,W)=(32,128,128)$. For ground images, initial feature maps of $(32,64,256)$ are extracted; these are then projected into BEV using our feature Gaussians projection, resulting in final ground BEV features also of dimensions $(32,128,128)$. Our model is trained using the AdamW optimizer [loshchilov2017decoupled] (weight decay $10^{-3}$) and a OneCycleLR scheduler [smith2019super] with a cosine annealing strategy, where the learning rate peaks at $6.25\times 10^{-5}$. We use a batch size of 8 on a single 4090 GPU, training for 8 epochs on KITTI and 14cf epochs on VIGOR.